Suppose that in a random selection of 100 colored candies, 27% of them are blue. The candy company claims that the percentageIdentify P-value for this hypothesis test is _.

( round to three decimals places)
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p =0.2

Ha : p \neq 0.2

\hat p = 0.27

P0 = 0.2

1 - P0 = 1-0.2=0.8

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.27-0.2/ [\sqrt(0.2*0.8) / 100]

= 1.75

P(z > 1.75) = 1 - P(z <1.75 ) = 1-0.9599=0.0401

P-value = 2*0.0401=0.0802

P VALUE 0.080